# Compression, Impact and Hot Rebound Flows from Coronal Rain Downflows

### v1

@@ 4.4 Rain Impact on the Lower Atmosphere
Figure 17: Rain impact detected by AIA and SJI. Left column: snapshot of AIA 193 Å  (top) and SJI 1400 Å  (bottom) immediately prior to impact, where the white cubic spline denotes the trajectory of the rain clump. Right column: the corresponding time-distance diagrams along the white cubic spline in AIA (top) and SJI (bottom). The time-distance diagrams clearly shows the impact of downflow 1 (cyan arrow) with smaller clumps impacting after (green arrow). The rebound flow can only be seen in the AIA 193 Å  time-distance diagram. See the online animation corresponding to this Figure.
@@
Figure 19: The top and bottom rows show average intensities for SJI 1400 Å  (left) and AIA 193 Å  (right) prior and during impact, respectively. The averages are done with the time-distance diagrams of Figure 17 over the time interval 30.531.530.5-31.5 min (prior to impact) and 333433-34 min (during impact).
Figure 20: Similar to Figure 19 but for HRIEUV. The average intensity was obtained with the time-distance (white spline) diagram of Figure 10 over the time interval 30.531.530.5-31.5 min (prior to impact) and 33.53433.5-34 min (during impact).
@@
To compare more closely the observed intensity variability across the channels we plot in Figure 19 a time average over 1 min prior (t=[30.531.5]t=[30.5-31.5]~min) and during the impact (t=[3334]t=[33-34]~min) for the intensities seen in the time-distance diagrams. Both the ‘steady’ intensity enhancement and that from the impact can be seen as very similar large bumps in the intensity profiles (with an increased value for the impact). We measure the widths of these enhacements by fitting a single or double Gaussian over the profile, as shown in the Figure. The SJI profiles prior and during the impact can be fitted with a single Gaussian with a FWHM of 0.98 Mm and 1.29 Mm, respectively. Similarly, AIA can be fitted with a double Gaussian, with a narrow (first) component’s FWHM of 1.18 Mm and 1.30 Mm, respectively, prior and during impact. The broad (second) component of the fit has a FWHM of 4.76 Mm and 4.25 Mm, respectively, prior and during impact. We repeat this measurement for HRIEUV using the time-distance diagram of Figure 10, and we show the resulting intensity profile and Gaussian fit in Figure 20. Also for this case, the intensity profile is best fitted by a double Gaussian. Prior to impact, the narrow (first) component has a FWHM of 0.58 Mm, and the broader (second) component has a FWHM of 3.833.83~Mm. In contrast with the AIA and SJI profiles, during the impact the HRIEUV intensity profile is more significantly affected. The widths of the narrow and broad components then become 2.0 Mm and 3.77 Mm. These changes could be due to the change in the LOS.
The widths of both the narrow and broad components for SJI 1400 Å  HRIEUV and AIA 193 Å  are very similar, implying that the heating (either steady, prior to impact) or impulsive (during impact), affects the entire transition region. The transition region is naturally narrower in SJI 1400 Å  compared to AIA 193 Å, but it is interesting that it is even narrower in HRIEUV, in agreement with the results of Mandal et al. (2023) where the studied dynamic fibrils with HRIEUV appear to show a bright front with a thickness of less than 1 Mm. The different LOS of HRIEUV could also affect the observed width. The fact that during impact, the AIA 193 Å  and SJI 1400Å  intensity enhancements (which share the same LOS) are observed at the same height in both channels suggests no change in the observed impact depth of the coronal rain clumps in SJI 1400 Å  compared to AIA 193 Å  further indicating that the entire transition region is affected by the impact. The persistent narrow bump observed in HRIEUV, and AIA 193 Å  (and in many other AIA channels) could be interpreted as a footpoint localised steady heating component, as expected by TNE theory. To have a measure of the width of this transition we measure the thickness from 10% the peak of the Gaussian fits to 10% the background value on the left side of the Gaussian fits (this is marked by the green points in each Panel of Figures 19) and 20. In the case of AIA and HRIEUV, we do this with the Gaussian fit on the small bump, since this is the one located lower in the atmosphere. We find values for SJI 1400 Å  of 0.61 Mm and 0.92 Mm prior and during impact, respectively. Similarly, for HRIEUV we find 1.14 Mm and 1.28 Mm. Finally, for AIA 193 Å  we find 1.71 and 1.77 Mm.
@@ 4.5 Rebound Flow
After the impact of downflow 1, a clear upward propagating feature originating from the impact region, termed ‘upward flow 1’, was identified between t=3437t=34-37 min in AIA 193 Å  and AIA 211 Å, respectively, in Figures 18 and 21. A second upward propagating feature can be seen (best in AIA 193 Å) between t=4448t=44-48 min (‘upward flow 2’), following the impact of downflow 2. Upward flow 1 has constant projected speed of $83.6\pm 10.0$~km s-1 in both AIA 193 Å and 211 Å. Similarly, upward flow 2 has a constant projected speed of $85.0\pm 10.0$~km s-1 in AIA 193 Å. Both rebound flows occur almost instantly following the impact of the descending coronal rain. Hence, we interpret these propagating features as rebound flows because of the relatively low speeds compared to the sound speed.
The fact that we observe these upflows very clearly in AIA 193 Å  and 211 Å  indicates relatively hot temperatures reaching 1.2521.25-2 MK. This is supported by the DEM analysis in Figure 14, which shows that only temperature bins corresponding to a temperature range of $\log(T)=6.15$ to 6.25 revealed the rebound flow. Upon close inspection, other wavelengths such as AIA 94 Å, 131 Å, 171 Å, and HRIEUV, also show signatures (but less distinct) of these upward flows denoted by the black arrows in the respective time-distance diagrams. The distance along the loop the upflows are seen to reach in both the AIA 193 Å  and 211 Å  is approximately $58\pm 3$ Mm. $53.3\pm 10$~km s-1 Using the projected and total velocity of downflow 1 and rearranging equation 7, we obtain $\theta_{AIA}=0.19\text{ radians}$. Using this, we can calculate the total velocity of upward flows 1 and 2 in Figure 18. Taking the projected velocity of upward flow 1 to be $83.6\,\text{km/s}$, we get a total velocity of $85.2\,\text{km/s}$. Taking the projected velocity of upward flow 2 to be $85.0\,\text{km/s}$, we get a total velocity of $v=86.5\,\text{km/s}$.
@@ 4.6 Energy Estimate
We obtain an electron number density for the clump of $(5.86\pm 0.37)\times 10^{11}$ cm-3, which is in line with usually reported density values for coronal rain (Antolin & Froment, 2022). Based on the rain volume (assuming cylindrical geometry), and using the morphological properties mentioned in section 4.2.1 and observed total descending speeds calculated in section 4.2.2, we obtain a kinetic energy for the rain clump of $(6.19\pm 2.04)\times 10^{25}$ erg.
We also calculate the thermal and kinetic energy of the rebound flow to analyse how much energy from the rain clump is transferred to it. The former is given by $U=\frac{3}{2}k_{B}VTn_{\text{tot}}$, where $V$, $T$ and $n$ denote, respectively, the volume, temperature and number density of the upflow. Using the values for the upward flow found in section 4.5 we find thermal and kinetic energies of $(1.01\pm 0.25)\times 10^{24}$ erg and $(9.24\pm 0.8)\times 10^{24}$ erg, respectively. We notice a factor of more than 10 difference between these values and the rain clump energy.
@@
We therefore find that less than $20\%$ of the clump’s kinetic energy is transferred to the rebound flow. Since the impact is seen across all EUV and UV wavelengths (except AIA 1600 and 1700) we can deduce that most of the energy is radiated away in the transition region upon impact. For convenience we summarise all calculated values in Table 1.
@@
Width (km) $549\pm 27$ $862\pm 7$
Length (km) $8702\pm 1200$ $7000\pm 500$
Kinetic Energy (erg) $(6.19\pm 2.0)\times 10^{25}$ $(1.01\pm 0.25)\times 10^{24}$
Thermal Energy (erg) $(4.69\pm 0.86)\times 10^{24}$ $(9.24\pm 0.8)\times 10^{24}$
@@ 5 Discussion and Conclusion
Throughout the entirety of the event, we observed strong EUV variability both during the rain shower (where downflows 1 and 2 are observed) and after its impact on the lower atmosphere. At the start of the observation, the coronal strands appeared densely packed and uniform in intensity, and by the end of the observation, we observed multiple strands disappearing with new ones appearing and notable gaps between strands. Such EUV variability is in agreement with that reported from loops in TNE-TI (Antolin et al., 2023), also in line with the generally observed variability in coronal loops (Ugarte-Urra et al., 2006; Antolin & Froment, 2022).
@@
Using the separation angle between AIA 171 Å  and HRIEUV at the time, we were able to calculate the total rain speed of downflows 1 and 2 to be $72.5\pm 10.0$ km s-1 and $86.4\pm 10.0$ km s-1, respectively. No clear acceleration or deceleration was observed, suggesting a balance of gas pressure and gravity. The speeds obtained are in agreement with previous studies (De Groof et al., 2005; Antolin & Rouppe van der Voort, 2012; Ishikawa et al., 2020). These dynamics are likely linked to the observed intensity brightening ahead of the rain suggesting compression. Compression from coronal rain (or prominence downflow) has been analysed analytically and numerically(Mackay & Galsgaard, 2001; Adrover-González et al., 2021), predicting deceleration or piston-like oscillations around nodal points along the loop. Oliver et al. (2014) have demonstrated that the abrupt formation of a condensation generates sounds waves upstream and downstream, that modify the pressure as they travel and effectively decelerates the clump and leads to a constant velocity during most of the downfall. Recently, this effect has been explained by Hillier et al. (2025) as a dynamic mass contribution from the compressed material ahead of the rain, which then modifies the momentum of the rain.
Based on the EUV absorption feature observed by HRIEUV, we calculated the width of downflow 1 to be $549\pm 27$ km. This width falls in line with those found with the same instrument at similar resolution by (Antolin et al., 2023), and off-limb with SJI 2796 Å, 1400 Å, and 1330 Å  at similar resolution with IRIS by (Antolin et al., 2015), strengthening the previous result that coronal rain widths are independent of temperature. As in Şahin et al. (2023), the width of of the clump was observed to be constant along the loop’s length. This property remains unexplained, and may be attributed to the fundamental scales over which coronal heating operates Antolin & Froment (2022) and/or to the pressure balance that follows thermal instability.
@@
We also detect upward motions that follow the impact of downflows 1 and 2. Rebound flows and shocks produced by the rain’s impact have been reported first by Antolin et al. (2023) with HRIEUV. In our case, the upward motions are only clearly observed in the AIA 193 Å  and 211 Å  passbands, indicating temperatures of $\log T=6.1-6.3$, supported by the DEM analysis. The upward motions travel at a constant total velocity of $(85.1-86.5)\pm 10.0$ km s-1. Since these speeds are much lower than the sound speed we interpret them as hot rebound flows, which then start to reheat the loop, indicating the start of a new TNE-TI cycle. Such atmospheric responses are expected from numerical simulations (e.g. Müller et al., 2003; Antolin et al., 2022). These upward motions can be traced clearly up to the formation height of downflow 1. The upflows are seen only about a minute after the downflow impact, further supporting the presence of multiple clumps accompanying downflows 1 and 2. The widths of upflow 1 is $\approx 860$~km, significantly larger than the rain width. This could be explained by the lower spatial resolution of AIA. However, an additional effect may contribute to these relatively large widths. It is plausible that the rain’s impact generates transverse magnetic perturbations in the low transition region, due to the higher plasma $\beta$ at those heights, which, in turn, should generate compression in the neighboring field lines, leading to upward flows in a wider volume than that traced by the rain. The future higher resolution observations in hot lines offered by MUSE and EUVST should be able to confirm this hypothesis.
@@
One of the most visible effects of the rain in the present case is its impact on the lower atmosphere. The rain shower, composed of small clumps, produces strong and rapid variability on the order of minutes observed in all AIA EUV channels, in HRIEUV and the SJI 1400 Å  channel. Reale et al. (2013) report prominence fallback from very large heights to produce similarly strong intensity brightenings across all EUV wavelengths. However, they also report emission in the AIA UV (low chromospheric) wavelengths, which are likely due to the higher falling speeds (up to 450km s-1, probably due to the negigible effect of the magnetic field in their case) and consequently larger kinetic energies. The strong brightenings we observe are accompanied by a more steady and lower intensity enhancement over tens of minutes, matching roughly the duration of the rain shower, visible in AIA 131 Å, 171 Å  and HRIEUV. This variation of intensity is likely due to the inhomogeneity present in the rain. The fact that the impact is seen in SJI 1400 Å  indicates that the condensations reach the low transition region. The width and location of the brightening caused by the impact is very similar between SJI 1400 Å, HRIEUV and AIA 193 Å, which suggests that the transition region thickness is very thin (below $\approx 1200$ km), matching predictions from numerical modelling (e.g. Reale, 2014). However, in the latter we observe a secondary width enhancement of larger magnitude, indicative of heating.
@@
We estimated the duration of downflow 1 to be $\approx 3$ min, with a kinetic energy of $(6.19\pm 2.0)\times 10^{25}$ erg. We also estimated the rain shower to have a duration of roughly 35 min. However, according to the time-distance diagrams (notably AIA 304 Å), we can see that the bulk of the rain shower lasts 15 min (where the EUV absorption is predominant in AIA 304 Å). Hence, if we assume a roughly similar density, the total energy of the rain shower is roughly $3.09\times 10^{26}$ erg, that is, on the order of a microflare. It is very likely that the rain’s structure is highly inhomogeneous. In the calculation of the rain’s density we assumed a temperature of 10,000 K for the rain (which is likely too high). Finally, the length of the rain is inferred rather than directly observed. In any case, these assumptions should compensate each other to some extent.

### v2

@@ 4.4 Rain Impact on the Lower Atmosphere
Figure 17: Rain impact detected by AIA and SJI. Left column: snapshot of AIA 193 Å  (top) and SJI 1400 Å  (bottom) immediately prior to impact, where the white cubic spline denotes the trajectory of the rain clump. Right column: the corresponding time-distance diagrams along the white cubic spline in AIA (top) and SJI (bottom). The time-distance diagrams clearly shows the compression produced by downflow 1 (horizontal cyan arrow), followed by the impact (vertical green arrow). The rebound flow can only be seen in the AIA 193 Å  time-distance diagram. See the online animation corresponding to this Figure.
@@
Figure 19: The average intensities for SJI 1400 Å (top), HRIEUV (middle) and AIA 193 Å (bottom), before (left) and during impact (right), respectively. The averages are done with the white spline of their respective time-distance diagrams (Figure 17 and 10) over the time interval 30.531.530.5-31.5 min (before impact) and 333433-34 min (during impact).
@@
To compare more closely the observed intensity variability across the SJI 1400 Å, HRIEUV and AIA 193 Å  channels we plot in Figure 19 a time average over 1 min prior (t=[30.531.5]t=[30.5-31.5]~min) and during the impact (t=[3334]t=[33-34]~min) for the intensities seen in the time-distance diagrams. In this comparison it is worth to have in mind that while IRIS and AIA share the same LOS, Solar Orbiter does not. However, since the LOS is not so different (angle between both instruments is $\approx 20^{\circ}$), we expect the changes produced by the different LOS to be small. This LOS difference introduces however an error in the co-alignment of features. Since we are not interested in an exact one-to-one correspondence at subarcsecond level (which is not possible given the different LOS), we address this uncertainty by taking the average over the time-distance diagrams for all the splines shown in Figures 4 and 5, which fully cover the rain impact.
@@
The panels in Figure 19 are organised in terms of temperature formation of the dominant spectral line in the passband, with lowest to largest from top to bottom. Both the ‘steady’ intensity enhancement and that from the impact can be seen as very similar large bumps in the intensity profiles (with an increased value for the impact). We measure the widths of these enhacements by fitting a single, double or tripple Gaussian over the profile, as shown in the Figure. While the peak increases by a factor of two during the impact, the shape of the SJI profile is not greatly affected by the impact, and can be fitted with a single Gaussian prior and during the impact, with a FWHM of 0.98 Mm and 1.29 Mm, respectively. For HRIEUV, the peak intensity also increases but only by a factor of 1.3. The HRIEUV intensity profile is more structured around the peak and is best fitted by a tripple Gaussian. Prior and during impact, the narrowest (first) component has a FWHM of 0.53 Mm and 0.61 Mm, respectively. Similarly, the second component has a FWHM of 2.74 Mm and 1.89 Mm, respectively. Lastly, the third and broadest component has a FWHM of 9.67 Mm and 6.88 Mm, respectively. For AIA, the increase due to impact is smaller in amplitude, with a factor of 1.1. Prior and during impact, AIA can be well fitted with a double Gaussian, with a narrow (first) component’s FWHM of 1.18 Mm and 1.30 Mm, respectively. The broad (second) component of the fit has a FWHM of 4.76 Mm and 4.25 Mm, respectively.
It is very likely that the narrow profiles (in the case of SJI 1400 Å  the only profile) correspond to the transition region. The fact that the intensity increase is largest in this channel suggests that the rain clump reaches at least the low transition region, where it probably deposits most of its energy. We note that the width of the SJI 1400 Å  profile is similar to that of the second component of HRIEUV and the narrow component of AIA 193 Å (prior and during impact). In contrast, the narrow width of HRIEUV is significantly smaller. These similarities and differences are likely due to differences in the response functions combined with differences in the spatial resolutions. For instance, SJI and HRIEUV have similar spatial resolution, but SJI 1400 Å  is sensitive to $10^{4.8}$~K plasma, while HRIEUV has a broader response with a peak at $10^{6}$~K and includes significant cooler contribution at $10^{5.5~}$K. Similarly, AIA has a factor of $\approx 4$ lower spatial resolution, and AIA 193 Å  peaks at $10^{6.15}$~K but also includes cool contribution from $10^{5.4}$~K onwards. This is supported by the time-distance diagram of the emission measure of Fig. 14, where we can see that the bright region near the footpoint for the emission measure bins $\log T=5.55-6.15$ is about half the size ($\lesssim 5$~Mm) of the corresponding bright region for the emission measure bins $\log T=6.15-6.25$ ($\lesssim 10$ Mm). We also note that the AIA and SJI profiles show little change in their widths and peak location during the impact, which suggests a very narrow transition region less than 1.2 Mm in thickness. On the other hand, the HRIEUV intensity profile is more significantly affected. This reflects the higher sensitivity of the upper transition region to energy injection at lower heights, in agreement with the results of Mandal et al. (2023) where the studied dynamic fibrils with HRIEUV appear to show a bright front with a thickness of less than 1 Mm.
Prior to the impact, we note that the broad (second and third components) observed in HRIEUV, and AIA 193 Å  (and in many other AIA channels) are persistent, which could be interpreted as a footpoint localised steady heating signature, as expected by TNE-TI theory. In this case, the FWHM of these components, which vary between 2 Mm and 10 Mm, could be a measure of the heating scale height.
@@ 4.5 Rebound Flow
After the impact of downflow 1, a clear upward propagating feature originating from the impact region, termed ‘upward flow 1’, was identified between t=3437t=34-37 min in AIA 193 Å  and AIA 211 Å, respectively, in Figures 18 and 20. A second upward propagating feature can be seen (best in AIA 193 Å) between t=4448t=44-48 min (‘upward flow 2’), following the impact of downflow 2. Upward flow 1 has constant projected speed of $83.6\pm 10.0$~km s-1 in both AIA 193 Å and 211 Å. Similarly, upward flow 2 has a constant projected speed of $85.0\pm 10.0$~km s-1 in AIA 193 Å. Both rebound flows occur almost instantly following the impact of the descending coronal rain. Hence, we interpret these propagating features as rebound flows because of the relatively low speeds compared to the sound speed.
The fact that we observe these upflows very clearly in AIA 193 Å  and 211 Å  indicates relatively hot temperatures reaching 1.2521.25-2 MK. This is supported by the DEM analysis in Figure 14, which shows that only temperature bins corresponding to a temperature range of $\log(T)=6.15$ to 6.25 revealed the rebound flow. Upon close inspection, other wavelengths such as AIA 94 Å, 131 Å, 171 Å, and HRIEUV, also show signatures (but less distinct) of these upward flows denoted by the black arrows in the respective time-distance diagrams. The distance along the loop the upflows are seen to reach in both the AIA 193 Å  and 211 Å  is approximately $58\pm 3$ Mm. $53.3\pm 10$~km s-1 Using the projected and total velocity of downflow 1 and rearranging equation 7, we obtain $\theta_{AIA}=0.19\text{ radians}$. Using this, we can calculate the total velocity of upward flows 1 and 2 in Figure 18. Taking the projected velocity of upward flow 1 to be $83.6\,\text{km~s}^{-1}$, we get a total velocity of $85.2\,\text{km~s}^{-1}$. Taking the projected velocity of upward flow 2 to be $85.0\,\text{km~s}^{-1}$, we get a total velocity of $v=86.5\,\text{km~s}^{-1}$.
@@ 4.6 Energy Estimate
We obtain an electron number density for the clump of $(5.86\pm 0.37)\times 10^{11}$ cm-3, which is in line with usually reported density values for coronal rain (Antolin and Froment, 2022). Based on the rain volume (assuming cylindrical geometry), and using the morphological properties (width of $0.55\pm 0.03$ Mm, length of $8.7\pm 0.3$ Mm) mentioned in section 4.2.1 and observed total descending speed (=72.5=72.5~km s-1) calculated in section 4.2.2, we obtain a kinetic energy for the rain clump of $(6.18\pm 1.86)\times 10^{25}$ erg. Using the temperature of 10,000 K for the rain clump we calculate the thermal energy, given by $U=\frac{3}{2}k_{B}VTn_{\text{tot}}$, where $V$, $T$ and $n_{\text{tot}}$ denote, respectively, the volume, temperature and number density of the rain clump. We find a value of $(4.79\pm 0.86)\times 10^{24}$ erg.
We also calculate the thermal and kinetic energy of the rebound flow to analyse how much energy from the rain clump is transferred to it. Using the values for the upward flow found in section 4.5 we find thermal and kinetic energies of $(1.01\pm 0.25)\times 10^{24}$ erg and $(9.24\pm 0.8)\times 10^{24}$ erg, respectively. We notice a factor of $\approx 6.5$ difference between these values and the rain clump energy.
@@
We therefore find that roughly $15\%$ of the clump’s kinetic energy is transferred to the rebound flow. Since the impact is seen across all EUV and UV wavelengths (except AIA 1600 and 1700) we can deduce that most of the energy is radiated away in the transition region upon impact. For convenience we summarise all calculated values in Table 1.
@@
Width (km) $549\pm 27$ $862\pm 7$ ${557}\pm 1$
Length (km) $8702\pm 300$ $7000\pm 500$
Kinetic Energy (erg) $(6.18\pm 1.86)\times 10^{25}$ $(1.01\pm 0.25)\times 10^{24}$
Thermal Energy (erg) $(4.79\pm 0.86)\times 10^{24}$ $(9.24\pm 0.8)\times 10^{24}$
@@ 5 Discussion and Conclusion
Throughout the entirety of the event, we observed strong EUV variability both during the rain shower (where downflows 1 and 2 are observed) and after its impact on the lower atmosphere. At the start of the observation, the coronal strands appeared densely packed and uniform in intensity, and by the end of the observation, we observed multiple strands disappearing with new ones appearing and notable gaps between strands. Such EUV variability is in agreement with that reported from loops in TNE-TI (Antolin et al., 2023), also in line with the generally observed variability in coronal loops (Ugarte-Urra et al., 2006; Antolin and Froment, 2022).
@@
Using the separation angle between AIA 171 Å  and HRIEUV at the time, we were able to calculate the total rain speed of downflows 1 and 2 to be $72.5\pm 10.0$ km s-1 and $86.4\pm 10.0$ km s-1, respectively. No clear acceleration or deceleration was observed, suggesting a balance of gas pressure and gravity. The speeds obtained are in agreement with previous studies (De Groof et al., 2005; Antolin and Rouppe van der Voort, 2012; Ishikawa et al., 2020). These dynamics are likely linked to the observed intensity brightening ahead of the rain suggesting compression. Compression from coronal rain (or prominence downflow) has been analysed analytically and numerically(Mackay and Galsgaard, 2001; Adrover-González et al., 2021), predicting deceleration or piston-like oscillations around nodal points along the loop. Oliver et al. (2014) have demonstrated that the abrupt formation of a condensation generates sounds waves upstream and downstream, that modify the pressure as they travel and effectively decelerates the clump and leads to a constant velocity during most of the downfall. Recently, this effect has been explained by Hillier et al. (2025) as a dynamic mass contribution from the compressed material ahead of the rain, which then modifies the momentum of the rain.
Based on the EUV absorption feature observed by HRIEUV, we calculated the width of downflow 1 to be $549\pm 27$ km. This width falls in line with those found with the same instrument at similar resolution by (Antolin et al., 2023), and off-limb with SJI 2796 Å, 1400 Å, and 1330 Å  at similar resolution with IRIS by (Antolin et al., 2015), strengthening the previous result that coronal rain widths are independent of temperature. As in Şahin et al. (2023), the width of of the clump was observed to be constant along the loop’s length. This property remains unexplained, and may be attributed to the fundamental scales over which coronal heating operates (Antolin and Froment, 2022) and/or to the pressure balance that follows thermal instability.
@@
Several works have shown that gas pressure ahead of the rain is the leading cause of its deceleration, and mostly constant velocity at the time of impact. Oliver et al. (2014); Martínez-Gómez et al. (2020) have shown that the terminal speeds depend on the density contrast between the rain clump and the loop’s density. Recently, Tan et al. (2023) in the context of cool clouds accreting towards galactic centres in the circumgalactic or intracluster medium, approximate the accretion braking process in the limit of small hydrodynamic drag as:
$v=\frac{m}{\dot{m}}(g_{\text{eff}}-\dot{v}),$ (14)
where we have modified their equation to account for the effective gravity $g_{\text{eff}}$ along a curved loop. Taking the average of $g_{\text{eff}}=0.171$ km s-2 for an elliptic loop with ratio of loop height to half baseline between 0.5 and 2, we have:
$t_{\text{grow}}=\frac{v_{CR}}{g_{\text{eff}}},$ (15)
where $v_{CR}=72.5$~km s-1 is the final velocity we find for downflow 1, and $t_{\text{grow}}$ is the growing time of the rain. We find, $t_{\text{grow}}\approx 7$~min, which corresponds well with the time over which the EUV absorption of downflow 1 is observed (see e.g. Fig. 12), and roughly half the time where we estimate downflow 1 to occur.
In our study, the compression produced by the rain is observed to be mostly isothermal while compressing the plasma to a maximum factor of 1.4. Based on the advective timescale being shorter than the conductive and radiative timescales, we assumed adiabatic conditions during the compression and calculated the ratio of specific heats to obtain a value of $\gamma=1.085\pm 0.218$. It is possible, however, that the compression just beneath the rain is unresolved, with a much stronger radiative region than currently observed. Also, from the accretion braking process we expect an important transfer of mass and momentum from the compressed region to the rain. Hillier et al. (2025) have shown that the numerical simulation results of rain downflows of Oliver et al. (2014); Martínez-Gómez et al. (2020) can be explained with the model of a virtual mass ahead of the rain. The energy of the compression would then go to slow down the cooling from the rain and perhaps reionise it as it falls, thereby explaining the isothermal behaviour of the compressed plasma. Hillier et al. (2025) obtain the following relation for the falling speed:
$v_{CR}=-\frac{g_{\text{eff}}H_{\text{eff}}/c_{s}}{1+\frac{2DH_{\text{eff}}}{F_{b}h_{b}}\sinh(\tau_{\text{eff}})}\tau_{\text{eff}},$ (16)
where we have also modified their equation to account for the effective gravity stated above, $H_{\text{eff}}$ is the effective pressure scale height for our loop at 1 MK, $F_{b}$ is the density contrast between the rain clump and the loop, $h_{b}$ is the length of the rain clump ahead of downflow 1, which we estimated to be $1.1\pm 0.5$ Mm, where the uncertainty includes the possibility of the rain being composed of smaller clumps. $D=0.6$ as given by Hillier et al. (2025), and $\tau_{\text{eff}}=c_{s}t/H_{\text{eff}}$ is the time during which downflow 1 occurs ($15\pm 2$ min). Replacing with our values we obtain total speeds of $v_{CR}=79.4\pm 17.7$ km s-1. In this calculation the main uncertainty is the length of the rain clump. Also, it is unclear how the assumed structuring (into smaller clumps) would affect on the dynamics. Despite the uncertainty, it is notable that the obtained rain speeds from accretion braking are very close to the observed total speed of the rain. Kohutova and Verwichte (2017) have conducted 2.5D-MHD simulations of coronal rain where, in addition to the gas pressure gradients, they show that magnetic pressure and tension can also play a role in decelerating the condensations. Density ratios (condensation to ambient) of 100 lead to piston-like effects where the condensations stop and bounce up and down prior to impact. High magnetic field strengths exacerbate this effect due to the bending of the field lines in the lower atmosphere. Only very high density ratios above 1000 (or even 10,000) effectively remove the piston effect. In our observations the density ratio is on the order of 170 and yet we do not observe this piston effect. The discrepancy may lie in the difference between 2.5D and 3D. Also, the boundary conditions implemented in the numerical models for the lower atmosphere can play a crucial role in the coronal rain dynamics.
Furthermore, we note that the faster speeds of downflow 2 are also consistent with accretion braking, since once the first clump falls the upstream material is more rarefied, leading to less compression and less transfer of mass and energy to the second clump (which is often observed in coronal rain). These results suggest that coronal rain can be used as a template for accretion braking, which is not possible in observations of the circumgalactic or intracluster media.
@@
We also detect upward motions that follow the impact of downflows 1 and 2. Rebound flows and shocks produced by the rain’s impact have been reported first by Antolin et al. (2023) with HRIEUV. In our case, the upward motions are only clearly observed in the AIA 193 Å  and 211 Å  passbands, indicating temperatures of $\log T=6.1-6.3$, supported by the DEM analysis. The upward motions travel at a constant total velocity of $(85.1-86.5)\pm 10.0$ km s-1. Since these speeds are much lower than the sound speed we interpret them as hot rebound flows, which then start to reheat the loop, indicating the start of a new TNE-TI cycle. Such atmospheric responses are expected from numerical simulations (e.g. Müller et al., 2003; Antolin et al., 2022), and have been shown to affect the subsequent thermal evolution of the loop (Johnston et al., 2025). These upward motions can be traced clearly up to the formation height of downflow 1. The upflows are seen only about a minute after the downflow impact, further supporting the presence of multiple clumps accompanying downflows 1 and 2. The widths of upflow 1 is $\approx 860$~km, more than 50% larger than the rain width. This could be explained by the lower spatial resolution of AIA. However, an additional effect may contribute to these relatively large widths. It is plausible that the rain’s impact generates transverse magnetic perturbations in the low transition region, due to the higher plasma $\beta$ at those heights, which, in turn, should generate compression in the neighboring field lines, leading to upward flows in a wider volume than that traced by the rain. The future higher resolution observations in hot lines offered by MUSE and EUVST should be able to confirm this hypothesis.
@@
One of the most visible effects of the rain in the present case is its impact on the lower atmosphere. The rain shower, composed of small clumps, produces strong and rapid variability on the order of minutes observed in all AIA EUV channels, in HRIEUV and the SJI 1400 Å  channel. Reale et al. (2013) report prominence fallback from very large heights to produce similarly strong intensity brightenings across all EUV wavelengths. However, they also report emission in the AIA UV (low chromospheric) wavelengths, which are likely due to the higher falling speeds (up to 450 km s-1, probably due to the negigible effect of the magnetic field in their case) and consequently larger kinetic energies. The strong brightenings we observe are accompanied by a more steady and lower intensity enhancement over tens of minutes, matching roughly the duration of the rain shower, visible in AIA 131 Å, 171 Å  and HRIEUV. This variation of intensity is likely due to the inhomogeneity present in the rain. The fact that the impact is seen in SJI 1400 Å  for which the intensity increase is the strongest (compared to HRIEUV and AIA 193 Å), indicates that the condensations reach the low transition region and deposit most of their energy there. The width and location of the brightening caused by the impact is very similar between SJI 1400 Å, HRIEUV and AIA 193 Å, which suggests that the transition region thickness is very thin (below $\approx 1200$ km), matching predictions from numerical modelling (e.g. Reale, 2014). During the impact, all hot components become narrower, which is expected from the heating.
We estimated the duration of downflow 1 to be $\Delta_{CR}\approx 3$ min, with a kinetic energy of $(6.18\pm 1.8)\times 10^{25}$ erg. The energy of the rain shower $K_{SH}$ can be written as:
$K_{SH}$ $=$ $K_{CR}N_{clumps}$ (17)
$\approx$ $K_{CR}\left(\frac{\Delta t_{SH}}{\Delta t_{CR}}\right),$ (18)
@@
where $N_{clumps}$ is the total number of rain clumps and $\Delta t_{SH}$ is the total duration of the rain shower, which we estimated to be roughly 35 min. In Eq. 18 we have made the reasonable assumption that the length of an individual clump times $N_{clumps}$ should be roughly equal to the length of the shower. According to the time-distance diagrams (notably AIA 304 Å), we can see that the bulk of the rain shower lasts 15 min (where the EUV absorption is predominant in AIA 304 Å). Hence, the total energy of the rain shower is roughly $4.64\times 10^{26}$ erg, that is, on the order of a microflare. This is likely a lower bound since in the calculation of the rain’s density we assumed a temperature of 10,000 K for the rain, which is likely too high. Finally, the length of the rain is inferred rather than directly observed. In any case, these assumptions should compensate each other to some extent.